OpenStudy (experimentx):

Integration:

OpenStudy (experimentx):

the answer of the integration is \[ {1 \over 4} \sqrt{b^2 -a^2}(b-a)\left( t + {\sin (4t)\over 4}\right)\]

OpenStudy (experimentx):

On the other hand, Mathematica expresses solution as \[ \frac{\sqrt{a-x} \sqrt{-b+x} \left(2 \sqrt{a-x} \sqrt{-b+x} (-a-b+2 x)-(a-b)^2 \text{ArcTan}\left[\frac{a+b-2 x}{2 \sqrt{a-x} \sqrt{-b+x}}\right]\right)}{8 \sqrt{(a-x) (-b+x)}} \]

OpenStudy (experimentx):

this is minus \[ {1 \over 4} \sqrt{b^2 -a^2}(b-a)\left( t - {\sin (4t)\over 4}\right) \]

OpenStudy (callisto):

Hmm.... \[\int \sqrt{(a-x)(x-b)}dx\]Let \(x=acos^2t + bsin^2t\) \[dx = (-2a\sin t\cos t + 2b \sin t\cos t)dt = (b-a)sin2t dt\] Then, the integral becomes (a-x) = a - (acos^2 t + bsin^2 t) = a (sin^2 t) - b sin^2 t = (a-b) sin^2 t (x-b) = (acos^2t + bsin^2t) - b = acos^2 t - b (cos^2 t) = (a-b) cos^2t \[\int \sqrt{(a-b)sin^2t \times (a-b)cos^2t} (b-a)sin2tdt\]\[=\int [(a-b)sint cost] (b-a)sin2tdt\]\[=-\frac{1}{2}\int (b-a)^2sin^22tdt\] \[=-\frac{1}{2}(b-a)^2\int sin^2 2tdt\]\[=-\frac{1}{2}(b-a)^2\int \frac{1-cos4t}{2}dt\]\[=-\frac{1}{4}(b-a)^2 (t-\frac{1}{4}sin4t)+C\]Anything wrong here?

OpenStudy (experimentx):

looks like i made error

OpenStudy (experimentx):

yeah i made error ... - ( 1 - cos ..) = -

OpenStudy (experimentx):

If we put \[ \sin (t) = \sqrt{x-a\over b-a} \\ \cos (t) = \sqrt{b-x\over b-a} \\ t = \arctan \sqrt { \left( x-a \over b-x\right) } \]

OpenStudy (experimentx):

we get pretty close to